# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v2

@@ B.1 Ablation Study of Scaling Factor
B.2 Ablation Study of ATPL and Co-Learnability
Table 3 breaks down zero-shot solved rates on 6 held-out BW tasks for TRACED and its ablations at 10k updates. TRACED achieves the top score in five of six tasks (Basic, Hardcore, Stairs, Stump, Roughness) and ranks second in the remaining one. The ATPL only variant (TRACED – CL) leads to the first place in one task (PitGap) and second place in two tasks (Basic, Stump), while the Co-Learnability only variant (TRACED – ATPL) only reaches the second place in two tasks (Hardcore, Roughness). Removing both components (the ACCEL baseline) ranks second on one (Stairs). These results confirm that both transition prediction error and Co-Learnability are essential to TRACED’s high and stable performance across diverse tasks.
Table 3: BipedalWalker Ablation Results. Mean $\pm$ SE of zero-shot test returns on 6 held-out BipedalWalker tasks, comparing ACCEL (10k), TRACED - ATPL (10k), TRACED - CL (10k), and TRACED (10k, $\alpha=1.5$, $\beta=0.6$) over 5 seeds. Bold indicates the best result per task; underline indicates the second-best.
@@ B.2 Ablation Study of ATPL and Co-Learnability
B.3 Hyperparameter Sensitivity Analysis
Table 4 reports the impact of varying the ATPL weight $\alpha$ and the Co-Learnability weight $\beta$ in TRACED on 12 held-out MiniGrid mazes, each evaluated after 10k PPO updates. Four of the six $(\alpha,\beta)$ settings outperform the 10k update baselines (Appendix E), and even the weakest setting matches the performance of PLR at 10k. This demonstrates that TRACED delivers consistently strong zero-shot generalization across a wide range of hyperparameters, without the need for extensive tuning.
Table 4: Effect of Hyperparameter Weights $(\alpha,\beta)$ on MiniGrid Generalization. Mean $\pm$ SE of zero-shot solved rates across 12 held-out MiniGrid environments, evaluated under various $(\alpha,\beta)$ settings. Except for the final column, TRACED with $(\alpha=1.0,\beta=1.0)$, averaged over five seeds, all other entries report single-seed measurements. Bold indicates the best result per task; underline indicates the second-best.
@@ B.3 Hyperparameter Sensitivity Analysis
Table 5 reports zero-shot test returns on six held-out BipedalWalker terrains after 10k PPO updates, under five different $(\alpha,\beta)$ settings. Notably, three of those configurations outperform all 10k-update baselines (Appendix E). Even the mean return across all five $(\alpha,\beta)$ choices ($\approx 64.9$) exceeds every baseline. These results demonstrate that TRACED delivers consistently strong zero-shot generalization across a wide range of weight settings, without requiring extensive hyperparameter tuning.
Table 5: Effect of Hyperparameter Weights $(\alpha,\beta)$ on BipedalWalker Generalization. Mean $\pm$ SE of zero-shot test returns on six held-out BipedalWalker terrains, for various $(\alpha,\beta)$ settings. Only the final column (TRACED with $(\alpha=1.5,\beta=0.6)$) is averaged over five seeds; all other entries are single-seed runs. Bold indicates the best result per terrain; underline indicates the second-best.
@@
Basic 282.06 291.04 281.32 279.37 293.67$\pm$3.56
Hardcore –7.60 115.87 29.55 50.10 86.83$\pm$17.96
Stairs –52.78 –21.94 –42.37 –66.02 –29.0$\pm$10.4
PitGap 12.96 –52.72 –52.21 –122.54 –39.26$\pm$11.42
Stump –100.95 –40.11 –74.83 –73.93 34.16$\pm$54.58
Roughness 148.22 187.43 244.57 185.83 193.29$\pm$21.6
Mean 46.99 79.93 64.34 42.13 89.95$\pm$12.95
Appendix C Efficiency Analysis
C.1 Wall-Clock Training Time in BipedalWalker
Table 6: Wall-Clock Training Time on BipedalWalker. Average wall-clock duration (hours $\pm$ SE over five runs) for each algorithm at 10k and 20k PPO updates.
Table 6 reports wall-clock training times on the BipedalWalker domain. At 10k updates, TRACED requires $35.64\pm 0.53$ h, about 6% more than ACCEL’s $33.26\pm 1.17$h, yet delivers a 22% increase in mean solved rate (see Section 4).
C.2 Sample Complexity: Environment Steps
Environment PPO Updates PLR ADD ACCEL TRACED
MiniGrid 10k 82M 41M 93M 93M
MiniGrid 20k 165M 82M 185M
BipedalWalker 10k 165M 80M 174M 174M
BipedalWalker 20k 329M 160M 347M
Appendix D Visualizing Curriculum Dynamics
@@
Figure 12: Representative terrains from BipedalWalker selected by ATPL and co-learnability at two training stages. The top two rows show terrains after 2.5k PPO updates (first row: low ATPL; second row: high ATPL), and the bottom two rows after 5k updates (third row: low ATPL; fourth row: high ATPL). Within each row, the left column displays low co-learnability examples and the right column high co-learnability examples.
@@
Figure 12 illustrates the joint impact of ATPL and Co-Learnability on terrain selection at two training milestones (2.5k and 5k PPO updates). ATPL alone captures the raw challenge level: low-ATPL terrains (first and third rows) remain relatively smooth, while high-ATPL terrains (second and fourth rows) feature larger gaps and steeper bumps. As training proceeds from 2.5k to 5k updates, the overall terrains become systematically harder, demonstrating TRACED’s ability to ramp up difficulty in lockstep with the agent’s improving skills. Co-Learnability then refines this progression by filtering out extremes: low Co-Learnability (left column) tends to generate either trivial or overwhelmingly difficult levels that can stall learning, whereas high Co-Learnability (right column) favors intermediate-difficulty terrains that transfer more effectively across tasks. Together, these visualizations confirm that ATPL drives a steadily increasing complexity trajectory, while Co-Learnability smooths it to sustain robust, transferable learning throughout the curriculum.
Appendix E Numerical Results
E.1 Minigrid Environment
Table 8: MiniGrid Zero-Shot Solved Rates. Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks for DR, PLR, ADD, ACCEL, and TRACED, evaluated after 10k and 20k PPO updates over 10 seeds. Bold indicates the best result per task; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
20k 0.51 $\pm$ 0.13 0.86 $\pm$ 0.06 0.64 $\pm$ 0.12 0.65 $\pm$ 0.1
20k 0.27 $\pm$ 0.11 0.55 $\pm$ 0.15 0.18 $\pm$ 0.1 0.58 $\pm$ 0.14
20k 0.73 $\pm$ 0.02 0.8 $\pm$ 0.03 0.74 $\pm$ 0.04 0.85 $\pm$ 0.02
20k 0.5 $\pm$ 0.02 0.53 $\pm$ 0.03 0.61 $\pm$ 0.03 0.49 $\pm$ 0.03
20k 0.74 $\pm$ 0.1 0.61 $\pm$ 0.11 0.74 $\pm$ 0.04 0.22 $\pm$ 0.13
20k 0.54 $\pm$ 0.13 0.32 $\pm$ 0.11 0.72 $\pm$ 0.06 0.25 $\pm$ 0.12
20k 0.58 $\pm$ 0.16 0.92 $\pm$ 0.06 0.9 $\pm$ 0.1 0.8 $\pm$ 0.13
20k 0.18 $\pm$ 0.12 0.64 $\pm$ 0.15 0.63 $\pm$ 0.14 0.78 $\pm$ 0.13
Maze 10k 0.02 $\pm$ 0.01 0.13 $\pm$ 0.1 0.04 $\pm$ 0.02 0.41 $\pm$ 0.13 0.7 $\pm$ 0.1
20k 0.1 $\pm$ 0.1 0.52 $\pm$ 0.16 0.2 $\pm$ 0.13 0.59 $\pm$ 0.14
20k 0.16 $\pm$ 0.11 0.56 $\pm$ 0.14 0.28 $\pm$ 0.1 0.65 $\pm$ 0.13
20k 0.46 $\pm$ 0.14 0.95 $\pm$ 0.04 0.54 $\pm$ 0.16 0.86 $\pm$ 0.09
20k 0.2 $\pm$ 0.05 0.55 $\pm$ 0.08 0.36 $\pm$ 0.08 0.53 $\pm$ 0.08
20k 0.41 $\pm$ 0.03 0.65 $\pm$ 0.04 0.55 $\pm$ 0.04 0.6 $\pm$ 0.05
Table 9: MiniGrid Ablation Results. Mean $\pm$ SE of zero-shot solved rates on 12 held-out MiniGrid tasks, comparing ACCEL (10k), TRACED – ATPL (10k), TRACED – CL (10k), and TRACED (10k) over 5 seeds. For TRACED and ACCEL, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation. Bold indicates the best result per task; underline indicates the second-best.
Environment ACCEL 10k TRACED - ATPL 10k TRACED - CL 10k TRACED 10k
16Rooms 0.7 $\pm$ 0.19 0.72 $\pm$ 0.14 0.91 $\pm$ 0.06 0.79 $\pm$ 0.19
16Rooms2 0.44 $\pm$ 0.17 0.01 $\pm$ 0.01 0.79 $\pm$ 0.17 0.72 $\pm$ 0.17
SimpleCrossing 0.77 $\pm$ 0.02 0.68 $\pm$ 0.04 0.86 $\pm$ 0.04 0.89 $\pm$ 0.01
FourRooms 0.46 $\pm$ 0.04 0.36 $\pm$ 0.02 0.48 $\pm$ 0.04 0.47 $\pm$ 0.02
SmallCorridor 0.03 $\pm$ 0.03 0.44 $\pm$ 0.21 0.57 $\pm$ 0.2 0.49 $\pm$ 0.17
LargeCorridor 0.11 $\pm$ 0.1 0.42 $\pm$ 0.22 0.65 $\pm$ 0.21 0.5 $\pm$ 0.14
Labyrinth 0.55 $\pm$ 0.23 0.4 $\pm$ 0.24 0.49 $\pm$ 0.2 1.0 $\pm$ 0.0
Labyrinth2 0.88 $\pm$ 0.1 0.16 $\pm$ 0.11 0.62 $\pm$ 0.18 0.98 $\pm$ 0.01
Maze 0.42 $\pm$ 0.24 0.29 $\pm$ 0.19 0.63 $\pm$ 0.2 0.59 $\pm$ 0.17
Maze2 0.42 $\pm$ 0.24 0.3 $\pm$ 0.19 0.41 $\pm$ 0.15 0.42 $\pm$ 0.19
Maze3 0.8 $\pm$ 0.2 0.68 $\pm$ 0.2 0.86 $\pm$ 0.12 0.86 $\pm$ 0.07
PerfectMaze(M) 0.33 $\pm$ 0.04 0.34 $\pm$ 0.02 0.53 $\pm$ 0.07 0.66 $\pm$ 0.11
Mean 0.49 $\pm$ 0.04 0.4 $\pm$ 0.04 0.65 $\pm$ 0.03 0.7 $\pm$ 0.04
Table 10: PerfectMaze Zero-Shot Solved Rates. Mean $\pm$ SE of zero-shot success rates on PerfectMazeLarge and PerfectMazeXL after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, and TRACED over 10 seeds. Bold indicates the best result per environment; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
20k 0.02 $\pm$ 0.01 0.19 $\pm$ 0.09 0.18 $\pm$ 0.1 0.20 $\pm$ 0.08
20k 0.0 $\pm$ 0.0 0.12 $\pm$ 0.09 0.14 $\pm$ 0.09 0.09 $\pm$ 0.06
Table 8 reports per-task zero-shot solved rates on the 12 held-out MiniGrid mazes, following [21]. We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (at both 10k and 20k). TRACED at 10k achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best rate in one maze (Maze2). Averaged across all 12 mazes, TRACED attains a mean solved rate of 0.66, outperforming every other baseline configuration.
Table 9 presents zero-shot solved rates on 12 held-out MiniGrid mazes for TRACED and its ablations at 10k updates. Full TRACED attains the highest score in six mazes (SimpleCrossing, Labyrinth, Labyrinth2, Maze3, PerfectMaze(M), and FourRooms) and ranks second in the other six, with a mean solved rate of 0.70, outperforming every variant. These results confirm that both transition prediction error and Co-Learnability are critical to TRACED’s superior, consistent performance across diverse tasks.
Table 10 compares zero-shot solved rates of TRACED against DR, PLR, ADD, and ACCEL on two large-scale PerfectMaze environments: PerfectMazeLarge and PerfectMazeXL. After 10k updates, TRACED achieves the highest success rate on PerfectMazeLarge (27% ± 7%) and the third-highest on PerfectMazeXL (10% ± 4%), matching ADD at 20k. These results underscore TRACED’s ability to scale to extreme maze sizes.
E.2 BipedalWalker Environment
Table 11 compares TRACED’s zero-shot returns against DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains. At 10k updates, TRACED attains the highest mean return in three environments (Basic, Hardcore, and Stump) and achieves an overall mean of 89.95 $\pm$ 31.72, outperforming all baselines.
Table 11: BipedalWalker Zero-Shot Test Returns. Mean $\pm$ SE of episode returns on six held-out BipedalWalkerHardcore terrains after 10k and 20k PPO updates, comparing DR, PLR, ADD, ACCEL, ACCEL-CENIE, and TRACED over 5 seeds. Bold indicates the best result per terrain; underline indicates the second-best. ACCEL-CENIE is estimated by applying the original CENIE paper’s ACCEL-to-ACCEL-CENIE performance ratio [20] to our ACCEL implementation.
Environment Update DR PLR ADD ACCEL ACCEL-CENIE TRACED
20k -1.55 $\pm$ 8.83 -1.89 $\pm$ 8.93 14.63 $\pm$ 21.99 62.69 $\pm$8.91 75.81
Table 12 compares zero-shot solved rates across six BipedalWalker terrains for DR, PLR, ADD, ACCEL, and TRACED. A trajectory is considered “solved” if its return exceeds 230. At 10k updates, TRACED achieves the top solved rate in five of the six tasks (Basic, Hardcore, Stairs, PitGap, Stump) and ranks second on Roughness, yielding an overall mean of 0.36—substantially higher than ACCEL’s 0.29 and the other baselines. No baseline matches TRACED’s combined efficiency and reliability in zero-shot generalization.
Table 12: BipedalWalker Zero-Shot Solved Rates. Mean $\pm$ SE of solved rates, defined as the fraction of episodes exceeding a return of 230, on 6 held-out BipedalWalkerHardcore terrains (Basic, Hardcore, Stairs, PitGap, Stump, Roughness) after 10k and 20k PPO updates. Results compare DR, PLR, ADD, ACCEL, and TRACED over 5 seeds. Bold indicates the best per terrain; underline indicates the second-best.
Environment Update DR PLR ADD ACCEL TRACED
20k 0.16 $\pm$ 0.13 0.35 $\pm$ 0.22 0.2 $\pm$ 0.2 0.99 $\pm$ 0.0
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.04 $\pm$ 0.04 0.23 $\pm$ 0.08
Stairs 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.02 0.03 $\pm$ 0.02
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.01
PitGap 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.01
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.02
Stump 10k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.18 $\pm$ 0.12
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.02 $\pm$ 0.0.2 0.0 $\pm$ 0.0
20k 0.0 $\pm$ 0.0 0.0 $\pm$ 0.0 0.03 $\pm$ 0.03 0.76 $\pm$ 0.03
20k 0.03 $\pm$ 0.02 0.06 $\pm$ 0.04 0.05 $\pm$ 0.05 0.34 $\pm$ 0.03
Appendix F Implementation Details
F.1 Minigrid Environment

### v3

@@ A.1 Ablation Study on the Scaling Factor
Deviating from defaults on either $\alpha$ or $\beta$ reduces the solved rates in most domains (Table 3). For both ignoring one of the weighted terms ($\alpha=0.0$ or $\beta=0.0$) and using excessively large weight ($\alpha=100.0$ or $\beta=100.0$), solved rates show slight improvement for some individual tasks, but statistical drops for overall performance. The default TRACED weights show the best performance compared to fixed projection baselines. This result shows that an effective balance between ATPL and Co-Learnability is needed.
A.2 Ablation Study on ATPL and Co-Learnability
In this ablation, we evaluate TRACED by removing each of its two components in turn. Table 4 reports zero-shot solve rates on six held-out BipedalWalker tasks after 10k PPO updates. Results are averaged over five random seeds and reported as mean $\pm$ standard error.
Table 4: Ablation study on ATPL and Co-Learnability. Comparing ACCEL (10k), TRACED - ATPL (10k), TRACED - CL (10k), and TRACED (10k, $\alpha=1.5$, $\beta=0.6$). Bold indicates the best result per task; underline indicates the second-best.
@@ A.2 Ablation Study on ATPL and Co-Learnability
TRACED achieves the top score in five of six tasks (Basic, Hardcore, Stairs, Stump, Roughness) and ranks second in the remaining one. The ATPL only variant (TRACED - CL) leads to the first place in one task (PitGap) and second place in two tasks (Basic, Stump), while the Co-Learnability only variant (TRACED - ATPL) only reaches the second place in two tasks (Hardcore, Roughness). Removing both components (the ACCEL baseline) ranks second on one (Stairs). These results confirm that both transition-prediction error and Co-Learnability are essential to TRACED’s high and stable performance across diverse tasks.
A.3 Hyperparameter Sensitivity Analysis
We analyze impact of varying the ATPL weight $\alpha$ and the Co-Learnability weight $\beta$ in TRACED across the 12 held-out MiniGrid environments and six held-out BipedalWalker terrains. Results are reported for a single random seed due to computational constraints.
Table 5: Effect of hyperparameter weights $(\alpha,\beta)$ on MiniGrid environments. Evaluations are performed under varying $(\alpha,\beta)$ settings. Bold indicates the best; underline indicates the second-best.
@@ A.3 Hyperparameter Sensitivity Analysis
In MiniGrid environments, four of the six $(\alpha,\beta)$ settings outperform the 10k-update baselines (Appendix P), and even the weakest setting matches the performance of PLR at 10k (Table 5). These results indicate that TRACED provides consistently strong zero-shot generalization across a broad range of hyperparameters, without extensive tuning.
Table 6: Effect of hyperparameter Weights $(\alpha,\beta)$ on BipedalWalker environments. Evaluations are performed under varying $(\alpha,\beta)$ settings. Bold indicates the best; underline indicates the second-best.
@@
[…]
In BipedalWalker, three of the configurations outperform all 10k-update baselines (Appendix P) (Table 6). Even the mean return across all five $(\alpha,\beta)$ choices $(\approx 64.9)$ exceeds every baseline. These results demonstrate that TRACED delivers consistently strong zero-shot generalization across a wide range of weight settings, without requiring extensive hyperparameter tuning.
A.4 Ablation Study on the Number of Workers
We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel. The only difference between our experimental setup and ACCEL (Parker-Holder et al., 2022) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively. To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers. Even under this setting, TRACED consistently outperforms ACCEL, indicating that our gains are robust to the PPO worker configuration (Table 7).
Table 7: Ablation study on Number of Workers. Bold indicates the best; underline indicates the second-best.
Environment ACCEL 5k ACCEL 10k TRACED 5k TRACED 10k
16Rooms $0.37\pm 0.17$ $0.78\pm 0.14$ $\textbf{0.95}\pm 0.02$ $\underline{0.93}\pm 0.07$
16Rooms2 $0.20\pm 0.20$ $\underline{0.77}\pm 0.23$ $0.29\pm 0.14$ $\textbf{0.91}\pm 0.04$
SimpleCrossing $0.65\pm 0.09$ $0.73\pm 0.07$ $\underline{0.78}\pm 0.02$ $\textbf{0.88}\pm 0.05$
FourRooms $0.33\pm 0.08$ $\underline{0.45}\pm 0.04$ $0.33\pm 0.04$ $\textbf{0.59}\pm 0.01$
SmallCorridor $\underline{0.52}\pm 0.21$ $0.27\pm 0.17$ $\textbf{0.62}\pm 0.31$ $0.10\pm 0.04$
LargeCorridor $\underline{0.42}\pm 0.17$ $0.36\pm 0.30$ $\textbf{0.49}\pm 0.24$ $0.16\pm 0.13$
Labyrinth $0.15\pm 0.15$ $0.49\pm 0.29$ $\underline{0.67}\pm 0.33$ $\textbf{0.91}\pm 0.09$
Labyrinth2 $0.0\pm 0.0$ $0.51\pm 0.29$ $\underline{0.65}\pm 0.20$ $\textbf{1.0}\pm 0.0$
Maze $0.33\pm 0.33$ $0.36\pm 0.32$ $\underline{0.95}\pm 0.04$ $\textbf{0.99}\pm 0.01$
Maze2 $0.01\pm 0.01$ $\textbf{0.74}\pm 0.23$ $0.22\pm 0.19$ $\underline{0.60}\pm 0.20$
Maze3 $0.14\pm 0.13$ $0.48\pm 0.27$ $\underline{0.63}\pm 0.29$ $\textbf{1.0}\pm 0.0$
PerfectMaze(M) $0.20\pm 0.09$ $0.48\pm 0.07$ $\underline{0.58}\pm 0.10$ $\textbf{0.63}\pm 0.13$
Mean $0.28\pm 0.10$ $0.54\pm 0.09$ $\underline{0.60}\pm 0.06$ $\textbf{0.72}\pm 0.09$
Appendix B Long-term Analysis on TRACED
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 12: TRACED Solved-Rate Time Series. Solved rate progression on MiniGrid tasks plotted over 0-45k PPO updates.
Maze
FourRooms
Labyrinth
PerfectMaze(M)
SimpleCrossing
SmallCorridor
Figure 13: ACCEL Solved-Rate Time Series. Solved rate progression on MiniGrid tasks plotted over 0-45k PPO updates.
To study long-horizon behavior, we ran single-seed, 45k PPO updates for both TRACED and ACCEL in the MiniGrid environment. Both methods exhibit post-convergence oscillations in success rate (i.e., solved rate). TRACED reaches near-peak success rapidly, but its performance subsequently oscillates rather than remaining perfectly stable (Figure 12). The same post-peak fluctuations appear under the ACCEL curriculum (Figure 13): although ACCEL attains its maximum more slowly, it shows a similar up-and-down pattern thereafter. These oscillations likely reflect inherent instability of the RL policy on held-out levels, and may be further influenced by ACCEL’s mutation dynamics, rather than a pathology introduced by our task-prioritization strategy. For this reason, we report TRACED’s performance at 10k updates (rather than 20k), before long-horizon oscillations confound comparisons. Mitigating these fluctuations in the student policy is an important direction for future work.
Appendix C Visualizing Curriculum Dynamics
@@
Figure 14: Representative terrains from BipedalWalker selected by ATPL and co-learnability at two training stages. The top two rows show terrains after 2.5k PPO updates, and the bottom two rows after 5k updates.
@@
Figure 14 illustrates the joint impact of ATPL and Co-Learnability on terrain selection at 2.5k and 5k PPO updates. ATPL alone captures the raw challenge level: low-ATPL terrains (first and third rows) remain relatively smooth, while high-ATPL terrains (second and fourth rows) feature larger gaps and steeper bumps. As training proceeds from 2.5k to 5k updates, the overall terrains become systematically harder, demonstrating TRACED’s ability to ramp up difficulty in lockstep with the agent’s improving skills. Co-Learnability then refines this progression by filtering out extremes: low Co-Learnability (left column) tends to generate either trivial or overwhelmingly difficult levels that can stall learning, whereas high Co-Learnability (right column) favors intermediate-difficulty terrains that transfer more effectively across tasks. Together, these visualizations confirm that ATPL drives a steadily increasing complexity trajectory, while Co-Learnability smooths it to sustain robust, transferable learning throughout the curriculum.
Appendix D Level Evolution
D.1 Visualization of Level Evolution in Minigrid
The visualization of how the Minigrid environment evolves as the number of blocks increases (Figure 15). Each step of the evolutionary process produces an edited level that has a high learning efficiency.
D.2 Evolution of Levels in BipedalWalker
Figure 16 shows the complexity metric results trained by three methods, TRACED (Ours), ACCEL + ATPL, ACCEL. Starting with plain terrain (near zero point), all three methods guide levels with more complex terrain. With aspect of (a) Stump height, (b) Stump height high, (e) Stair height step metrics, the results show TRACED and ACCEL + ATPL quickly evolve levels compared to ACCEL. This rapid level increasing brings, short wall-clock relative to performance and high performance results in various test environments.
(a) Stump height low
(b) Stump height high
(c) Stair height low
(d) Stair height high
(e) Stair height step
(f) Pit gap low
(g) Pit gap high
(h) Roughness
Figure 16: Emergent BipedalWalker terrain complexity metrics. Aspect of (a) Stump height low, (b) Stump height high, and (e) Stair height step, complexity grows faster under TRACED (Ours, pink) or ACCEL + ATPL (brown) than under ACCEL (blue). The result is averaged over five random seeds. This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
D.3 Visualization of Level Evolution in BipedalWalker
Figure 17 shows the visualization of how the environment evolves as the complexity of the task increases. The process starts from a plain terrain and gradually evolves to a level of terrain with increasingly complex parameters.
Appendix E Agent Trajectory Visualizations Across Environments
E.1 MiniGrid
E.2 BipedalWalker
Appendix F Efficiency Analysis
F.1 Wall-Clock Training Time in BipedalWalker
Table 8 reports wall-clock training times on the BipedalWalker domain, in hours (mean $\pm$ s.e.) over five runs at 10k and 20k PPO updates. At 10k updates, TRACED requires $35.64\pm 0.53$ h, about 7% more than ACCEL’s $33.26\pm 1.17$ h, yet achieves a 22% higher mean solved rate (Section 4).
Table 8: Wall-Clock Training Time on BipedalWalker. Average wall-clock duration for each algorithm
F.2 Sample Complexity: Environment Steps
Environment PPO Updates PLR ADD ACCEL TRACED
MiniGrid 10k 82M 41M 93M 93M
MiniGrid 20k 165M 82M 185M
BipedalWalker 10k 165M 80M 174M 174M
BipedalWalker 20k 329M 160M 347M
Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. We followed ACCEL (Parker-Holder et al., 2022) in all respects except the procedure used to compute task priority. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
If the buffer is full, we evict the lowest‐priority task before appending. Otherwise, with probability $p_{\mathrm{replay}}$ the teacher enters the replay+mutation phase, drawing a batch of $B$ tasks from the buffer proportional to their TaskPriority scores, updating the policy on each, and recomputing regret. We then select the $N_{\mathrm{mutate}}$ tasks with the smallest regret, mutate each via the editor, evaluate the new variant, and replace its parent in the buffer. Finally, for every modified task we recompute TaskDifficulty, CoLearnability, and TaskPriority (Eq. 7) before proceeding to the next timestep.
Algorithm 1 UED Workflow with Transition‐Prediction Loss and Co‑Learnability
3:while policy has not converged do
4:  Sample decision $d_{t}\sim\mathrm{Bernoulli}(p_{\mathrm{replay}})$
[…]
6:   Sample new task $\theta\sim\mathcal{G}$
8:   Compute regret estimate
9:   Append $\theta$ to buffer $\Lambda$ (if $|\Lambda|>K$, remove lowest‐priority)
[…]
11:   Sample a batch $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ w.p. $\propto$ $\frac{1}{\mathrm{TaskPriority}(\cdot,t)}$
12:   for $k=1\ldots B$ do
13:     Collect trajectory $\tau_{k}$ by executing $\pi_{\phi}$ on $\theta_{k}$
14:     Update policy $\phi$ using rewards from $\tau_{k}$
15:     Compute $\widehat{\mathrm{Regret}}(\tau_{k})$
16:   end for
18:   for each selected $\theta_{k^{\prime}}$ do
19:     $\tilde{\theta}\leftarrow$ editor($\theta_{k^{\prime}}$)
20:     Collect trajectory $\tilde{\tau}$ on $\tilde{\theta}$ (no gradient)
[…]
22:     Replace $\theta_{k^{\prime}}$ in $\Lambda$ with $\tilde{\theta}$
23:   end for
24:  end if
25:  Buffer update:
26:  for each task $i\in\Lambda$ updated at $t$ do
27:   Recompute $\mathrm{TaskDifficulty}(i,t)$
28:   Recompute $\mathrm{CoLearnability}(i,t-1)$
29:   Update $\mathrm{TaskPriority}(i,t)$ via Eq. 7
30:  end for
31:  $t\leftarrow t+1$
32:end while
Appendix H Implementation Details
H.1 Minigrid Environment